Circular No. 6881

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE The Supernova Cosmology Project (cf. IAUC 6804), reports the dis-

covery of 13 supernovae and 7 apparent supernovae, all found with the Cerro Tololo 4-m telescope in their High Redshift Supernova Search:

SN  1998 UT  α_{2000}  δ_{2000}  R  z  type
1998as Mar. 22 10^{h}53^{m}56.^{s}22 − 3^{o}39′10′′.0 22.0 0.36E Ia
1998at Mar. 22 10 54 54.37 − 3 44 10.0 23.8 0.40 II
1998au Mar. 22 10 55 46.11 − 4 19 51.0 23.5  ?
1998av Mar. 22 10 56 07.98 − 4 11 20.6 23.3  ?
1998aw Mar. 22 10 56 15.02 − 4 15 21.5 25.0 0.44* Ia
1998ax Mar. 22 10 57 16.38 − 3 33 59.2 22.3 0.50E Ia
1998ay Mar. 22 10 57 21.51 − 3 14 49.9 22.9 0.64 Ia
1998az Mar. 22 10 57 21.78 − 3 25 23.0 22.7 0.46 Ia
1998ba Mar. 22 13 43 36.92 + 2 19 30.6 22.3 0.43E Ia
1998bb Mar. 22 13 43 53.58 + 1 25 33.7 24.0 0.57 Ia
1998bc Mar. 22 13 45 10.14 + 2 23 52.4 22.1 0.40E Ia
1998bd Mar. 22 13 45 29.05 + 1 57 32.5 21.2 0.25E Ia
1998be Mar. 22 13 46 20.89 + 2 02 40.4 23.8 0.64 Ia
1998bf Mar. 22 13 46 31.80 + 2 08 37.7 24.3  ?
1998bg Mar. 22 13 47 11.74 + 2 07 22.4 20.9  ?
1998bh Mar. 22 13 47 36.55 + 2 18 22.5 23.1  ?
1998bi Mar. 22 13 47 44.72 + 2 20 57.2 23.7 0.75 Ia
1998bj Mar. 22 13 47 49.63 + 1 59 47.8 23.4  ?
1998bk Mar. 22 13 48 48.25 + 2 58 41.2 22.5 0.41* Ia
1998bl Mar. 22 13 49 36.04 + 1 57 46.6 23.2  ?

The supernova spectra were obtained at the Keck II telescope on Mar. 30– Apr. 1, except for those marked ‘E’, which were obtained at the European Southern Observatory 3.6-m telescope on Mar. 28. The redshifts were determined from host-galaxy emission and absorption lines, except for those marked with an asterisk, which were determined solely from the spectra of the supernovae. The apparent supernovae were all followed with Cerro Tololo photometry on at least three nights. All supernovae were found to be within 2′′.5 of the host-galaxy center. R- and I-band photometry of the lightcurves are being obtained from the ground for all supernovae with determined redshifts. R-, I-, and J-band photometry for SNe 1998as, 1998aw, 1998ax, 1998ay, 1998ba, 1998bb, 1998be, and 1998bi are being obtained with the Hubble Space Telescope.

1998 April 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green